About how much smaller is the area of a square whose side measures 5 feet than the area of a circle whose diameter measures 6 feet?
Answer by ptaylor(2198)


Area of square=s^2=5^2=25 sq ft
Area of circle=pi*r^2=pi*(6/2)^2=9pi sq ft=9*3.14=28.26 sq ft
Area of circle - area of square=28.26-25=3.26 sq ft, so,
Area of square is 3.26 sq ft smaller

Does this help?------------ptaylor